OpenStudy (anonymous):

Describe how the graph of the function y=1.5sin(4x) is related to the graph of y=sin x.

OpenStudy (campbell_st):

well the standard form is y = asin(bx) a= amplitude of the curve amplitude is the max height of the curve and \[period = \frac{2\pi}{b}\] period is the distance along the horizontal axis needed to complete 1 cycle of the curve.

OpenStudy (anonymous):

Okay thank you, yeah I knew how to find the amplitude, period, phase shift, but I just didn't know how to set up the answer. Some teachers use standard form and some use a sentence format.

OpenStudy (anonymous):

There is no phase shift correct?

OpenStudy (campbell_st):

thats correct... well one way to do this is graph the curve, and then look at what's happening... the key is 4x so 4 cycles over a distance of 2pi

OpenStudy (anonymous):

Thanks for all the help campbell_st!